OpenStudy (deadshot):

Solve for x: 5 over quantity x squared minus 1 minus 2 over x equals 2 over quantity x plus 1?

OpenStudy (deadshot):

\[\frac{ 5 }{ x^2-1 }-\frac{ 1-2 }{ x }=\frac{ 2 }{ x+1 }\]

OpenStudy (anonymous):

same as last 1

OpenStudy (anonymous):

its 1-2?

OpenStudy (deadshot):

I think so, it was put as a word problem, but I think that's the equation

OpenStudy (deadshot):

oh, oops, no, it's \[\frac{ 5 }{ x^2-1 }-\frac{ 2 }{ x }=\frac{ 2 }{ x+1 }\]

OpenStudy (radar):